OpenStudy (anonymous):

A rectangular field will be fenced on all four sides. Fencing for the north and south sides costs $6 per foot and fencing for the other two sides costs $10 per foot. What is the maximum area that can be enclosed for $5400?

OpenStudy (anonymous):

i think something is wrong it is not right

OpenStudy (mathstudent55):

Let one dimension be called L The other dimension is called W The area of a rectangle is: A = LW The Perimeter, P = 2L + 2W The cost of the fencing is: 6*2L + 10*2W = 5400 12L + 20W = 5400 20W = 5400 - 12L W = 270 - 0.6L Substitute the expression for W into the expression for the Area A = L(270 - 0.6L) A = 270L - 0.6L^2 Take the derivative dA/dL = 270 - 1.2L Set the derivative equal to zero, and solve for L 270 - 1.2L = 0 1.2L = 270 L = 225 Substitute to get W W = 270 - 0.6(225) = 135 Length = 225 ft, Width = 135 ft A = 225 ft * 135 ft = 30,375 sq ft

OpenStudy (anonymous):

thanks God bless u it is right now. have great night.